OpenStudy (anonymous):

f(x) where f(x)if x+6≤4 and if 4x-6>4 as x approaches 4

OpenStudy (anonymous):

\[\lim (x ^{2}-2x-48)\div(x-8) \] as x approaches 8

OpenStudy (anonymous):

\[\lim_{x \rightarrow 8} \frac{x^2-2x-48}{x-8}=\lim_{x \rightarrow 8}\frac{\cancel{(x-8)}(x+6)}{\cancel{(x-8)}}\]